A car has a top speed of 71 m/s (about 159 mi/h) and has wheels of radius 35 cm. If the car starts from rest and reaches its operating speed after 7.1 s, what is the angular acceleration of the wheels? Assume the car accelerates uniformly.
rad/s2

explain please
Answer #1

linear acceleration = 71/7.1 = 10 m/s^2

angular acceleration = a/r = 10/0.35 = 28.57 rad/s^2
Answer #2

v1 = 0

w1 = v1/r1 = 0

v2 = 71 m/s

w2 = v2/r

= 71/0.35

= 202.86 rad/s


alfa = (w2-w1)/t

= (202.86-0)/7.1

= 28.6 rad/s^2
Answer #3

Radius of wheel, r=35cm=0.35m

Angular speed, w0=Linear speed/r=71/0.35=202.86 rad/sec

Fianl angular speed=0 rad/sec

Angular acceleration=[w-w0]/t

                                                =[202.86-0]/7.1

                                                =28.57 rad/s2